It takes one man one day to dig to day a 2m*2m*2m hole.How long does it takes 3 men
working at the same rate to dig a 4m*4m*4m hole ?

Answer by vleith(2983)

one man can dig 2*2*2 = 8metre cubed hole in a day.
How long would it take for that one guy to dig a hole 4*4*4 = 64 metre cubed?
64/8 = 8. It would take one guy 8 days.
You are told there are three guys, each able to dig at the same rate as the first guy.
8/3 = 2 2/3 days